OpenStudy (gabby1214):

Solve w+17=19 the solution is w=?

OpenStudy (anonymous):

w+17=19 Isolate w on one side. In order to do this, we need to subtract 17 (since it is being added, we need to do the opposite) from both sides. w+17=19 -17 -17 w = 2 <--- Answer. (:

OpenStudy (gabby1214):

Thank you for the help and explaining it!

OpenStudy (anonymous):

No problem!